# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Full Static Quantization

### v1

@@ 4.1 Overall Results
4.2 Latency and Memory Analysis
We measure the efficiency of MQuant in terms of speedup and memory usage under varying image resolutions, as well as the decode-stage acceleration introduced by AIFS. Unless otherwise noted, we adopt the common “text-image-text” input format (Duan et al., 2024) with 15 textual tokens, while the image size changes from $280\times 280$ to $5600\times 5600$.
Speedup and Memory Savings.
Overall Speedup As shown in Table 3, MQuant surpasses PyTorch BF16 and AWQ (W4-only) across all image resolutions, achieving up to 24.76% speedup over PyTorch at $840\times 840$. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability. In addition, our method provides memory savings over both PyTorch and AWQ, with reductions exceeding 100% compared to PyTorch (e.g., 152.92% at $840^{2}$). These benefits primarily arise from (1) eliminating the overhead of token-wise scale computation, and (2) converting mixed tokens into modality-decoupled tokens, avoiding slicing and concatenation when image resolutions become large. Decode-Stage Acceleration. We also measure decode time for generating 2,000 tokens with a custom W4A8 GEMV kernel (see Table5). Compared to per-token dynamic quantization, our AIFS+MSQ framework gains 23% speedup in the prefill stage and 100% speedup in decode time. By shifting online dynamic quantization to an offline static approach, we greatly reduce inference overhead, for long-sequence tasks. Moreover, since visual tokens are generally pricier than textual ones (Duan et al., 2024), these improvements translate to notable real-world cost reductions (e.g., $\approx 30\%$ for OpenAI-token-based pricing).
Table 6: Latency comparison under multi-turns setting.
Turns Config in a Turn All(s) Improve $\uparrow$
@@ 4.2 Latency and Memory Analysis
Table 7: The accuracy and speedup of our MSQ and AIFS scheme on the linear layer during prefill stage. The input sequence is structured as "text-image-text-image-text-image-text" with an image resolution of $2240\times 2240$ and 50 textual tokens. Latency were tested on an NVIDIA RTX 6000 Ada Generation.
[…]
BF16 BF16 84.43 93.87 842 2319 1.690 -
W4-g128(AWQ) BF16 83.93 (-0.50) 93.13 (-0.74) 828 (-14) 2252 (-67) 2.057 (+0.367) -17.8%
Acceleration for Multi-batch and Multi-turn Inference.
For multi-batch scenarios, we employ a “text-image-text” input setting at 2240$\times$2240 resolution, with each batch containing text lengths varying from 50 to 200, and uniformly generating 512 tokens during decoding. We use left-padding (pad_token_id) to align sequences of different lengths, masking out the padded regions to ensure they do not affect attention computations (Figure A.4 provides a schematic). Notably, our AIFS procedure remains fully compatible with multi-batch inputs, requiring no additional overhead. As shown in Table 4, MQuant consistently achieves around 20% faster inference across both prefill and decode stages compared to the floating-point baseline when the batch size ranges from 1 to 4. For multi-turn dialogue, we retain the same “text-image-text” format and resolution, with each turn containing 50 text tokens and a uniform decode length of 512 tokens. By maintaining key-value caches and position IDs across turns, we preserve context during the dialogue. As Table 6 shows, MQuant reduces end-to-end inference time by up to 38.7% over the floating-point baseline for 1–3 turns, showing the efficiency of our quantization approach in multi-turn settings.
4.3 Ablation Study
In this section, we select Qwen2-VL-7B (Wang et al., 2024), currently one of the most powerful open-source MLLMs, to ablate the effectiveness of our proposed designs.
Table 8: Ablation study of proposed quantization designs on Qwen2-VL-7B (Wang et al., 2024) with W4A8 setting.
[…]
Static AIFS + MSQ RMS
BF16 84.43 93.87 842 2319 6523
71.44 83.96 670 1911 5479
78.95 87.55 721 2095 5484
84.32 93.58 824 2255 5471
Ablation Study of Proposed Quantization Methods.
We perform ablations on Qwen2-VL-7B (Table 8) to isolate the contributions of each quantization component. Beginning with a baseline that applies GPTQ plus Hadamard transformations to both the LLM and vision parts, we progressively introduce MSQ and AIFS for the LLM. This significantly boosts accuracy, highlighting their effectiveness on text tokens. Next, leveraging offline and online Hadamard rotations in the vision encoder further improves results. Finally, incorporating RMS to suppress newly arising outliers achieves near-floating-point performance, underscoring the overall robustness and efficiency of our pipeline.
Figure 5: The Speedup of AIFS+MSQ on Qwen2-VL-7B.
Accuracy & Speedup via AIFS.
[…]
@@ 4.3 Ablation Study
Weight-only Quantization.
We compare our weight-only quantization results with the official Qwen2-VL-7B, which uses GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) to quantize only the LLM while keeping the visual encoder in BF16. For fairness, we adopt the same group size of 128. As shown in Table 6, our W4 LLM quantization aligns with their settings, achieving nearly lossless accuracy compared to other methods. Extending W4 quantization to the visual encoder enables 4-bit quantization of the entire MLLM; MQuant still performs comparably to the BF16 model and even surpasses other methods. Under the W4A8 weight-activation quantization setting, our results remain consistent, with some metrics surpassing advanced weight-only methods. These experiments confirm MQuant’s effectiveness and robustness across various quantization configurations, whether for weight-only or weight-activation quantization, and for partial or full quantization of MLLMs.
5 Conclusion
In this paper, we propose MQuant, an accurate and efficient post-training quantization (PTQ) framework specifically designed for MLLMs. Our approach addresses the unique challenges of MLLMs by applying Modality-Specific Static Quantization (MSQ) to handle distribution mismatches between vision and text tokens, and introducing an Attention-Invariant Flexible Switching (AIFS) mechanism to reduce time-to-first-token (TTFT) with static per-tensor scaling. Furthermore, we reveal that online Hadamard rotations can introduce severe weight outliers, and propose Rotation Magnitude Suppression (RMS) to effectively mitigate this issue. Extensive experiments on five mainstream MLLMs demonstrate that MQuant achieves near-floating-point accuracy under the W4A8 setting, highlighting its potential to advance MLLMs quantization and facilitate practical deployment in resource-constrained environments.
@@ Appendix A Appendix
A.1 MQuant Algorithm
Here, we present our MQuant algorithm for MLLMs in Algorithm 1.
@@ A.1 MQuant Algorithm
    Reorder the input mixed token sequence from $E={e^{t}_{1},\dots,e^{v}_{m},\dots,e^{v}_{n},\dots,e^{t}_{L}}$ to a unified modality-decoupled sequence $E^{u}={e^{v}_{m},\dots,e^{v}_{n},e^{t}_{1},\dots,e^{t}_{m-1},\dots,e^{t}_{n+1},\dots,e^{t}_{L}}$ using the proposed Attention-Invariant Flexible Switching (AIFS) scheme in Sec  3.1.
7:  Transform all the LayerNorm to RMSNorm in MLLM vision encoder $\bm{E}$ and Visual-Language Projector $\bm{P}$ using the proposed Post LayerNorm-to-RMSNorm transformation in Sec 3.2.
@@
12:  Address the weight outliers using caused by online Hadamard based on the proposed Rotation Magnitude Suppression (RMS) in Sec  3.2.
A.2 Rotary Position Embedding for Attention-Invariant Flexible Switching
@@
$R_{\Theta,m}^{d_{h}}=\begin{pmatrix}\cos{i\theta_{1}}&-\sin{i\theta_{1}}&0&0&\cdots&0&0\\ \sin{i\theta_{1}}&\cos{i\theta_{1}}&0&0&\cdots&0&0\\ 0&0&\cos{i\theta_{2}}&-\sin{i\theta_{2}}&\cdots&0&0\\ 0&0&\sin{i\theta_{2}}&\cos{i\theta_{2}}&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&\cos{i\theta_{d_{h}/2}}&-\sin{i\theta_{d_{h}/2}}\\ 0&0&0&0&\cdots&\sin{i\theta_{d_{h}/2}}&\cos{i\theta_{d_{h}/2}}\end{pmatrix}$ (9)
@@
In the proposed Attention-Invariant Flexible Switching (AIFS) mechanism, we also apply the rearrangement for position embedding to maintain the computation equivalence. For a mixed input token $E=\left\{e^{t}_{1},...,e^{v}_{m},...,e^{v}_{n},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$ (in Eq 3), where $m$ and $n$ denote the start and end indices of the visual tokens. Specifically, after AIFS, the unified token can formulated as: $E=\left\{e^{v}_{m},...,e^{v}_{n},e^{t}_{1},...,e^{t}_{m-1},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$. Therefore, the unified token indices after AIFS can be represented as:
$(m,...,n,1,...,{m-1},{n+1},...,L)=AIFS{(1,...,m,...,n,...,L)}$ (10)
@@
A.3 Weights Outliers after Online Hadamard Transformation
@@
$\bm{H}_{2}=\tfrac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]\qquad\textrm{and}\qquad\bm{H}_{2^{n}}=\bm{H}_{2}\otimes\bm{H}_{2^{n-1}}\,.$ (13)
$(AB)_{ij}=\sum_{r=1}^{n}a_{ir}b_{rj}=a_{i1}b_{1j}+a_{i2}b_{2j}+\cdots+a_{in}b_{nj}\vspace{-4mm}$ (14)
@@
A.4 Attention Mask in AIFS when Multi-batch Inference.
Figure 6: The illustration of causal mask when batch size > 1.
@@
A.5 Effectiveness of Rotational Magnitude Suppression for Weight Outliers in LLMs.
Table 10: WikiText-2 perplexity (PPL) on 4-bit quantized LLaMA2 models with an input sequence length of 2048. Lower PPL is better.
@@
For LLMs, compared to the original Quarot, integrating RMS with Quarot leads to performance improvements across LLaMA2 models with 7B, 13B, and 70B parameters, as detailed in Table 10. For MLLMs, ablation studies presented in Table 8 demonstrate that the RMS method significantly enhances quantization performance.
Figure 7: The number of prefill visual tokens across different MLLMs as the image splits or resolution increases.
A.6 Image Tokens in Various MLLMs
As shown in Fig 7, for different MLLMs (Bai et al., 2023b; Wang et al., 2024; Yao et al., 2024; Chen et al., 2024b), the number of prefill visual tokens grows as image resolution increases. This rapid expansion in token count exacerbates the inference latency, particularly in per-token dynamic quantization, which requires separate processing for each token, such as memory access and scale computation. As a result, the TTFT in MLLMs increases drastically, severely impacting overall inference latency. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating the increase in TTFT.
A.7 RMS Algorithm
@@
1:  Initialize an empty list $marks$
2:  for each linear layer in the $model$ do
3:    if Layer satisfies Equation 8 then
4:       Mark the layer and append its ID to $marks$
@@
7:  Apply Quarot or other Hadamard-based transformations to the $model$
8:  for each layer ID in $marks$ do
@@
11:  Quantize the $model$
12:  return $model$
A.8 Speedup and Memory Savings with Scaling Image Resolution
Table 11: Comparison of latency and memory saving with Pytorch and AWQ on Qwen2-VL-7B. means the Qwen2-VL-7B official implementation.
@@ A.8 Speedup and Memory Savings with Scaling Image Resolution
We fixed the input sequence as "text-image-text" with 15 textual tokens and presente the detailed changes of speed and memory, varying the image resolution from $280\times 280$ to $5600\times 5600$. Notably, the "text-image-text" sequence setting is not arbitrarily chosen; instead, it is a common setting in existing evaluation datasets (Duan et al., 2024). We evaluate speedup and memory savings by comparing PyTorch’s BF16, AWQ (W4-only), and our MQuant (W4A8). Speedup: As shown in Table 11, MQuant consistently achieves speedups over both PyTorch and AWQ across all resolutions, with a maximum of 24.76% over PyTorch at $840\times 840$. Notably, MQuant outperforms AWQ, which is slower than PyTorch at most resolutions due to negative speedups. This significant speedup highlights the advantage of our per-tensor static quantization, eliminating the overhead of token-wise scale computation. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability across various image sizes. Memory Savings: MQuant offers substantial memory reductions compared to both PyTorch and AWQ. It consistently reduces memory usage by over 100% compared to PyTorch (e.g., 152.92% at $840^{2}$) and significantly outperforms AWQ’s memory efficiency, achieving up to 101.07% savings at higher resolutions. These experiments demonstrate MQuant’s strengths in both latency and memory savings, achieving up to 24.76% faster inference and reducing memory consumption by over 100% compared to baseline methods. This makes MQuant a more efficient solution for deploying MLLMs in resource-constrained environments.
A.9 Comparison of Online vs. Without Online Hadamard Transform in Quarot

### v2

@@ 4.4. Latency and Memory Analysis
Inference Speedup and Memory Savings
Overall Speedup As shown in Table 11, MQuant surpasses BF16 and AWQ (W4-only) across all image resolutions, achieving up to 24.76% speedup over PyTorch at $840\times 840$. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability. Memory Efficiency. Our method provides memory savings over both BF16 and AWQ, with reductions exceeding 100% compared to PyTorch (e.g., 152.92% at $840^{2}$). These benefits primarily arise from (1) eliminating the overhead of token-wise scale computation, and (2) converting mixed tokens into modality-decoupled tokens, avoiding slicing and concatenation when image resolutions become large. More resolution experiments are in Appendix  A.8. Decoding Acceleration. We also measure decode time for generating 2,000 tokens with a custom W4A8 GEMV kernel (Table 12). Compared to per-token dynamic quantization, our AIFS+MSQ framework gains 23% speedup in the prefill stage and 100% speedup in decode stage. By shifting online per-toekn dynamic quantization to an offline per-tensor static approach, we greatly reduce inference overhead, especially for long-sequence tasks. Moreover, since visual tokens are generally pricier than textual ones (Duan et al., 2024), these improvements translate to notable real-world cost reductions (e.g., $\approx 30\%$ for OpenAI token pricing).
Table 13. Multi-Batch speedup comparison of MSQ + AIFS on W4A8 setting. Each row shows the cumulative total of text tokens, images, and textual responses for multi-turn inference.
[…]
Text Image Text bfp16 MQuant bfp16 MQuant bfp16 MQuant
1 10 $\times2240$ 50 2.54 1.93 +31.6% 18.01 12.89 +39.7% 20.55 14.82 +38.7%
2 10/10 $\times2240$ / $\times2240$ 50/100 5.42 4.15 +30.6% 37.82 31.56 +19.8% 43.24 35.71 +21.1%
3 10/10/10 $\times2240$ / $\times2240$ / $\times2240$ 50/100/150 8.24 6.42 +28.3% 48.03 40.35 +19.0% 56.27 46.77 +20.3%
4 10/10/10/10 $\times2240$ / $\times2240$ / $\times2240$ / $\times2240$ 50/100/150/200 11.17 8.67 +28.9% 59.09 49.92 +8.4% 70.26 58.59 +20.0%
Table 14. Latency comparison under multi-turns setting.
Turns Config in one Turn All(s) Improve $\uparrow$
@@
Acceleration for Multi-batch and Multi-turn Inference
For multi-batch scenarios, we use $\times2240$ resolution images with ”text-image-text” inputs, we evaluate batches containing 50-200 text tokens while generating 512 decoded tokens. Sequence alignment employs left-padding (pad_token_id) with proper masking (see Appendix A.4). AIFS maintains full multi-batch compatibility without overhead, enabling consistent 20% faster inference versus FP baselines across batch sizes 1-4 (Table 13). For multi-turn dialogue, under identical resolution and input format (50 text tokens per turn + 512 decoded tokens), MQuant achieves 38.7% faster end-to-end inference than BF16 for 1-3 turns (Table 14) through optimized key-value caches and position IDs across turns.
Table 15. Comparison under weight-only settings on Qwen2-VL-7B (Wang et al., 2024a). $\dagger$} means re-implementation based on the official weight-only setting with a group size of 128.
@@
Weight-only Quantization
We evaluate our weight-only quantization using GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) under consistent settings (group size=128), quantizing only the LLM while keeping the visual encoder in BF16. As shown in Table  15, our W4 LLM quantization achieves near-lossless accuracy comparable to existing methods. When extended to W4 quantization, MQuant maintains comparable performance with the BF16 baseline. The W4A8 quantization similarly shows competitive or superior results compared to weight-only approaches. These experiments validate MQuant’s effectiveness across: (1) weight-only vs. weight-activation schemes, and (2) partial vs. full-model quantization scenarios.
5. Conclusion
In this paper, we propose MQuant, an accurate and efficient PTQ framework specifically designed for MLLMs. Our approach addresses the challenges of MLLMs by applying Modality-Specific Static Quantization (MSQ) to handle distribution mismatches between visual and text tokens, and introducing an Attention-Invariant Flexible Switching (AIFS) mechanism to reduce TTFT with static per-tensor scaling. Furthermore, we reveal that online Hadamard rotations can lead to weight outliers, and propose Rotation Magnitude Suppression (RMS) to mitigate this issue. Extensive experiments on 5 mainstream MLLMs demonstrate that MQuant achieves near-floating-point accuracy under the W4A8 setting, highlighting its potential to advance MLLMs quantization and practical deployment in resource-constrained environments. For the quantization of video-based MLLMs, we also discuss in Appendix A.16.
@@ Appendix A Appendix
A.1. MQuant Algorithm
Here, we present overall MQuant algorithm for MLLMs in Algorithm 1.
@@ A.1. MQuant Algorithm
    Reorder the input mixed token sequence from $E={e^{t}_{1},\dots,e^{v}_{m},\dots,e^{v}_{n},\dots,e^{t}_{L}}$ to a unified modality-decoupled sequence $E^{u}={e^{v}_{m},\dots,e^{v}_{n},e^{t}_{1},\dots,e^{t}_{m-1},\dots,e^{t}_{n+1},\dots,e^{t}_{L}}$ using the proposed Attention-Invariant Flexible Switching (AIFS) scheme in Sec 3.1.
7: Transform all the LayerNorm to RMSNorm in MLLM vision encoder $\bm{E}$ and Visual-Language Projector $\bm{P}$ using the proposed Post LayerNorm-to-RMSNorm transformation.
@@
12: Address the weight outliers using caused by online Hadamard based on the proposed Rotation Magnitude Suppression (RMS) in Sec 3.2.
A.2. Rotary Position Embedding for Attention-Invariant Flexible Switching
@@
(10) $R_{\Theta,m}^{d_{h}}=\begin{pmatrix}\cos{i\theta_{1}}&-\sin{i\theta_{1}}&\cdots&0&0\\ \sin{i\theta_{1}}&\cos{i\theta_{1}}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&\cos{i\theta_{d_{h}/2}}&-\sin{i\theta_{d_{h}/2}}\\ 0&0&\cdots&\sin{i\theta_{d_{h}/2}}&\cos{i\theta_{d_{h}/2}}\end{pmatrix}$
@@
In the proposed Attention-Invariant Flexible Switching (AIFS) mechanism, we also apply the rearrangement for position embedding to maintain the computation equivalence. For a mixed input token $E=\left\{e^{t}_{1},...,e^{v}_{m},...,e^{v}_{n},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$ (in Eq. 3), where $m$ and $n$ denote the start and end indices of the visual tokens. Specifically, after AIFS, the unified token can formulated as: $E=\left\{e^{v}_{m},...,e^{v}_{n},e^{t}_{1},...,e^{t}_{m-1},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$. Therefore, the unified token indices after AIFS can be represented as:
(11) (m,,n,1,,m1,n+1,,L)=AIFS(1,,m,,n,,L)(m,...,n,1,...,{m-1},{n+1},...,L)=AIFS{(1,...,m,...,n,...,L)}
@@
A.3. Weights Outliers after Online Hadamard Transformation
@@
(14) $\bm{H}_{2}=\tfrac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]\qquad\textrm{and}\qquad\bm{H}_{2^{n}}=\bm{H}_{2}\otimes\bm{H}_{2^{n-1}}\,.$
(15) $(AB)_{ij}=\sum_{r=1}^{n}a_{ir}b_{rj}=a_{i1}b_{1j}+a_{i2}b_{2j}+\cdots+a_{in}b_{nj}\vskip-11.38109pt$
@@
A.4. Attention Mask in AIFS when Multi-batch Inference.
Figure 8. The illustration of causal mask when batch size > 1.
@@
A.5. Effectiveness of Rotational Magnitude Suppression for Weight Outliers in LLMs.
Table 16. WikiText-2 perplexity (PPL) on 4-bit quantized LLaMA2 models with an input sequence length of 2048. Lower PPL is better.
@@
For LLMs, compared to the original Quarot, integrating RMS with Quarot leads to performance improvements across LLaMA2 models with 7B, 13B, and 70B parameters, as detailed in Table. 4. For MLLMs, ablation studies presented in Table 7 demonstrate that the RMS method significantly enhances quantization performance.
Figure 9. The number of prefill visual tokens across different MLLMs as the image splits or resolution increases.
A.6. Image Tokens in Various MLLMs
As shown in Fig 9, for different MLLMs (Bai et al., 2023b; Wang et al., 2024a; Yao et al., 2024; Chen et al., 2024c), the number of prefill visual tokens grows as image resolution increases. This rapid expansion in token count exacerbates the inference latency, particularly in per-token dynamic quantization, which requires separate processing for each token, such as memory access and scale computation. As a result, the TTFT in MLLMs increases drastically, severely impacting overall inference latency. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating the increase in TTFT.
A.7. RMS Algorithm
@@
1:  Initialize an empty list marksmarks
2:  for each linear layer in the modelmodel do
3:   if Layer satisfies Eq. 9 then
4:    Mark the layer and append its ID to marksmarks
@@
7:  Apply Quarot or other Hadamard-based transformations to the modelmodel
8:  for each layer ID in marksmarks do
@@
11:  Quantize the modelmodel
12:  return modelmodel
Table 17. Comparison of latency and memory saving with Pytorch and AWQ on Qwen2-VL-7B. means the Qwen2-VL-7B official implementation.
@@
A.8. Speedup and Memory Savings with Scaling Image Resolution
We fixed the input sequence as "text-image-text" with 15 textual tokens and presented the detailed changes of speed and memory, varying the image resolution from $280\times 280$ to $5600\times 5600$. Notably, the "text-image-text" sequence setting is not arbitrarily chosen; instead, it is a common setting in existing evaluation datasets (Duan et al., 2024). We evaluate speedup and memory savings by comparing PyTorch’s BF16, AWQ (W4-only), and our MQuant (W4A8). Speedup: As shown in Table 17, MQuant consistently achieves speedups over both PyTorch and AWQ across all resolutions, with a maximum of 24.76% over PyTorch at $840\times 840$. Notably, MQuant outperforms AWQ, which is slower than PyTorch at most resolutions due to negative speedups. This significant speedup highlights the advantage of our per-tensor static quantization, eliminating the overhead of token-wise scale computation. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability across various image sizes. Memory Savings: MQuant offers substantial memory reductions compared to both PyTorch and AWQ. It consistently reduces memory usage by over 100% compared to PyTorch (e.g., 152.92% at $840^{2}$) and significantly outperforms AWQ’s memory efficiency, achieving up to 101.07% savings at higher resolutions. These experiments demonstrate MQuant’s strengths in both latency and memory savings, achieving up to 24.76% faster inference and reducing memory consumption by over 100% compared to baseline methods. This makes MQuant a more efficient solution for deploying MLLMs in resource-constrained environments.
A.9. Comparison of Online vs. Without Online Hadamard Transform in Quarot